# Bi-CoG: Bi-Consistency-Guided Self-Training for Vision-Language Models

### v1

@@ Introduction
To evaluate the robustness of Bi-CoG in open-world scenarios, we conduct a comprehensive study by integrating it with three representative prompt tuning methods across 14 datasets. For each dataset, we report two metrics: HM of base and novel class accuracies, and overall accuracy (Acc.). We set the base-to-novel class ratio to 50:50 and report the mean and standard deviation over three runs. As shown in Table 1, Bi-CoG consistently improves both metrics across all datasets, achieving the highest average performance. Notably, when applied to CoOp, Bi-CoG yields over a 5% improvement in HM. Moreover, Bi-CoG demonstrates particularly strong gains on traditional open-world benchmarks. For example, on low-resolution datasets such as CIFAR-10 and CIFAR-100 (Krizhevsky, Hinton et al. 2009), it achieves up to an 8% increase in HM, underscoring its effectiveness across datasets with varying resolutions. These results highlight the effectiveness of Bi-CoG and emphasize the importance of integrating unsupervised strategies with pre-trained models for improved generalization in open-world settings.
@@
As shown in Table 2, SSL algorithms exhibit a significant advantage over baseline prompt tuning methods on SSL datasets. This is largely due to the lower resolution of these datasets, which makes it more difficult for VLMs to effectively capture visual content (Krizhevsky, Hinton et al. 2009; Lv et al. 2025b). This limitation is also reflected in the sensitivity to the number of shots: increasing the number of shots from 4 to 25 yields much larger performance gains for SSL algorithms compared to VLM-based prompt tuning. However, Bi-CoG provides the model with more reliable training examples, enabling prompt tuning methods to achieve SOTA performance under the challenging 4-shot setting.
These results demonstrate that Bi-CoG can effectively leverage unlabeled data to capture richer visual representations, even in low-resolution and data-scarce scenarios.
@@
To further validate the effectiveness of Bi-CoG, we conducted comparative experiments against the current SOTA methods GRIP (Menghini, Delworth, and Bach 2023) and CPL (Zhang et al. 2024) on the Flowers102 dataset (Nilsback and Zisserman 2008). Following GRIP, we adopt the same base-to-novel class ratio of 62:38, and use CoOp as the underlying prompt tuning method for Bi-CoG and other approaches to ensure faireness. As shown in Table 3, our Bi-CoG achieves the best performance under both settings, outperforming CPL by up to 5.9 percentage points in SSL setting. More importantly, Bi-CoG eliminates the need for manually setting hyperparameters such as the number of pseudo-labels $k$, effectively decreasing the reliance on manual tuning and susceptibility to internal model bias.
In summary, Bi-CoG represents one of the most effective integrations of pretrain-finetuning and SSL methods to date.
Ablation Study
@@
Figure 3: (a) Pseudo-label accuracy over training iterations on the DTD dataset under different pseudo-labeling strategies; (b) Distribution of pseudo-labels on the EuroSAT dataset before and after applying the intra-model consistency.
@@
Table 3: Performance comparison on the Flowers102 dataset, with a base-to-novel class ratio of 62:38. Our method outperforms prior approaches under both settings.
Figure 4: Comparison between the estimated and true error rate ratios on StanfordCars under the open-world setting. The scaling factor is fixed at $\alpha=1$ throughout training.
@@
Table 4: Ablation study of Bi-CoG’s pseudo-labeling strategy under the open-world generalization setting. Experimental results show that each component plays a critical role in improving model performance.
To further validate the theoretical foundation of our error-aware mechanism, we analyze the evolution of the actual error rate ratio on pseudo-labeled samples and its estimation from labeled data on the StanfordCars dataset (Krause et al. 2013). As illustrated in Figure 4, the condition in Equation 8 holds reliably during training, supporting the practical validity of our theoretical derivation and confirming the soundness of the Bi-CoG’ dynamic self-training framework.
Finally, we conduct a comprehensive ablation study on the key components of Bi-CoG, using CoOp as the underlying prompt tuning method and reporting the average performance across 14 datasets. As shown in Table 4, any component of our three-stage pseudo-labeling strategy, which integrats bi-consistency and error-aware filtering, plays a critical role in improving model performance. This highlights the deeply coupled and effective design of the pseudo-labeling mechanism in Bi-CoG.
Conclusion
This paper investigates the integration of pre-trained vision-language models with semi-supervised learning for downstream tasks with sparse labeled data. We identify several critical issues in existing approaches, including the trade-off between label quality and model bias, as well as the limited adaptability arising from the reliance on manually set fixed thresholds. To address these challenges, we propose Bi-CoG, a plug-and-play self-training framework that incorporates a three-stage, deeply coupled pseudo-labeling strategy with adaptive adjustment capabilities. Theoretical analysis and experimental results across 14 datasets in both open-world generalization and semi-supervised settings demonstrate that Bi-CoG can be seamlessly integrated with any pretrain-finetuning methods, consistently achieving performance improvements.
Supplementary Material
A. Detailed Theoretical Analysis
A.1. Proof of Lemma 1
@@
When the aggregated training data contains noisy samples, let $\mathcal{F}$ denote the hypothesis space, $\eta$ the noise ratio in the training set ($\eta<0.5$), $\epsilon>0$ the target error rate, and $\delta<1$ the confidence parameter. According to PAC learning theory (Angluin and Laird 1988; Zhou and Li 2005), if the number of training samples $m$ satisfies
@@ A.2. Proof of Theorem 1
ImagetNet-100 Natural picture 100 130,000 5,000
@@ B. Implementation Details
We evaluate our method on 14 recognition datasets, comprising 11 commonly used benchmarks for evaluating vision-language models (VLMs) and 3 classical benchmarks for semi-supervised learning. These datasets include CIFAR-10 and CIFAR-100 (Krizhevsky, Hinton et al. 2009), ImageNet and ImageNet-100 (Deng et al. 2009), Caltech101 (Fei-Fei, Fergus, and Perona 2004), OxfordPets (Parkhi et al. 2012), StanfordCars (Krause et al. 2013), Food101 (Bossard, Guillaumin, and Van Gool 2014), Flowers102 (Nilsback and Zisserman 2008), FGVCAircraft (Maji et al. 2023), SUN397 (Xiao et al. 2010), UCF101 (Soomro, Zamir, and Shah 2012), DTD (Cimpoi et al. 2014), and EuroSAT (Helber et al. 2019).
For ImageNet, Caltech101, OxfordPets, StanfordCars, Flowers102, FGVCAircraft, SUN397, UCF101, DTD, and EuroSAT, we follow the dataset splits used in CoOp (Zhou et al. 2022a). For CIFAR-10, CIFAR-100, and ImageNet-100, we adopt the standard semi-supervised splits used in FixMatch (Sohn et al. 2020). Detailed statistics of these datasets are provided in Table 5. Specifically, for the open-world generalization experiment, we follow the standard base-to-novel setting(Zhou et al. 2022a), where base and novel classes are split evenly. We sample 16 shots per class as labeled data, and treat all remaining training samples as unlabeled. Due to computational constraints, for ImageNet we randomly sample 5% of the unlabeled set for training.
@@
As a plug-and-play method, Bi-CoG can be seamlessly integrated into any pretrain–finetune paradigm. In our experiments, we adopt CoOp (Zhou et al. 2022b), MaPLe (Khattak et al. 2023a), and PromptSRC (Khattak et al. 2023b) as the base methods, keeping all hyperparameters identical to those reported in their original papers. Unless otherwise specified, we set the number of models to $K=3$ and the scaling factor to $\alpha=1$. In the dynamic self-training stage, each model is first warmed up for $\zeta$ epochs using a small set of labeled data. Then, we apply our three-stage pseudo-label selection strategy to obtain pseudo-labels with high accuracy and low model bias, which are further used for self-training over $\kappa$ additional epochs. The detailed parameters are provided in Table 7. During the intra-model consistency stage, we apply weak and strong augmentations to each sample, with the specific augmentation strategies summarized in Table 6. We conduct all experiments on a single NVIDIA A800 GPU.

### v3

@@ 4.2 Experimental Results
To evaluate the robustness of Bi-CoG in open-world scenarios, we conduct a comprehensive study by integrating it with three representative prompt tuning methods across 14 datasets. Notably, in all experiments related to Bi-CoG, we controlled the random seeds during both the initialization and training phases for different models to ensure diversity in the initial training stages. For each dataset, we report two metrics: HM of base and novel class accuracies, and overall accuracy (Acc.). We set the base-to-novel class ratio to 50:50 and report the mean and standard deviation over three runs. As shown in Table 1, Bi-CoG consistently improves both metrics across all datasets, achieving the highest average performance. Notably, when applied to CoOp, Bi-CoG yields over a 5% improvement in HM. Moreover, Bi-CoG demonstrates particularly strong gains on traditional open-world benchmarks. For example, on low-resolution datasets such as CIFAR-10 and CIFAR-100 Krizhevsky et al. (2009), it achieves up to an 8% increase in HM, underscoring its effectiveness across datasets with varying resolutions. These results highlight the effectiveness of Bi-CoG and emphasize the importance of integrating unsupervised strategies with pre-trained models in open-world settings.
@@
As shown in Table 2, SSL algorithms exhibit a significant advantage over baseline prompt tuning methods on SSL datasets. This is largely due to the lower resolution of these datasets, which makes it more difficult for VLMs to effectively capture visual content Krizhevsky et al. (2009); Lv et al. (2025b). Notably, this advantage also comes at the cost of substantial computational overhead: conventional SSL methods typically require over 10 hours of training for the CIFAR dataset, whereas Bi-CoG achieves comparable or superior performance with only 1 hour of training under the same hardware configuration. This limitation of high computational cost, coupled with the sensitivity to the number of shots (increasing the number of shots from 4 to 25 yields much larger performance gains for SSL algorithms compared to VLM-based prompt tuning), highlights the practical challenges of traditional SSL methods. In contrast, Bi-CoG provides the model with more reliable training examples via adaptive pseudo-labeling, enabling prompt tuning methods to achieve SOTA performance under the challenging 4-shot setting while maintaining remarkable computational efficiency.
These results demonstrate that Bi-CoG can effectively leverage unlabeled data to capture richer visual representations even in low-resolution and data-scarce scenarios, while also significantly reducing training time compared to SSL methods, making it more practical and scalable.
@@
To further validate the effectiveness of Bi-CoG, we conducted comparative experiments against two SOTA methods: GRIP Menghini et al. (2023) and CPL Zhang et al. (2024). Specifically, GRIP defines a fixed epoch-dependent expression for selecting the number of pseudo-labels, while CPL relies on manually set confidence thresholds to filter pseudo-labels iteratively. Following the experimental protocol of GRIP, we adopted the same base-to-novel class ratio (62:38) and used CoOp as the underlying prompt tuning framework for Bi-CoG and all compared methods. This ensures a fair comparison by eliminating discrepancies from different base models or tuning strategies. As shown in Table 3, Bi-CoG achieves the best performance under both SSL and open-world settings, outperforming CPL by up to 5.9 percentage points. Notably, this superiority stems from Bi-CoG’s adaptive pseudo-label selection strategy during iterations: unlike GRIP’s epoch-dependent rule or CPL’s manual threshold tuning, Bi-CoG adjusts pseudo-label selection based on model consistency and estimated error rate, ensuring high-quality supervision signals. More importantly, Bi-CoG dispenses with the requirement for manually tuning hyperparameters, effectively reducing dependence on human expertise. In summary, Bi-CoG represents one of the most effective integrations of pretrain-finetuning and semi-supervised learning methods to date.
4.3 Ablation Study
@@
Figure 3: (a) Pseudo-label accuracy on DTD under different pseudo-labeling strategies; (b) Distribution of pseudo-labels on EuroSAT before and after applying the intra-model consistency.
Figure 4: Comparison between the estimated and true error rate ratios. The scaling factor is fixed at $\alpha=1$ throughout training.
@@
Table 3: Performance comparison on the Flowers102 dataset.
@@
Table 4: Ablation study of Bi-CoG’s pseudo-labeling strategy.
To further validate the theoretical foundation of our error-aware mechanism, we analyze the evolution of the actual error rate ratio on pseudo-labeled samples and its estimation from labeled data on StanfordCars Krause et al. (2013) under the open-world generalization setting. As illustrated in Figure 4, the condition in Equation 6 holds reliably during training, supporting the practical validity of our theoretical derivation and confirming the soundness of Bi-CoG.
Finally, we conduct a comprehensive ablation study on key components of Bi-CoG, using CoOp as the base model and reporting the average performance across 14 datasets. As shown in Table 4, any component of our three-stage pseudo-labeling strategy, which integrates bi-consistency and error-aware filtering, plays a critical role in improving model performance. This highlights the deeply coupled and effective design of the pseudo-labeling mechanism in Bi-CoG.
5 Conclusion
This paper investigates the integration of pre-trained VLMs with semi-supervised learning for downstream tasks with sparse labeled data. We identify critical issues in existing approaches, including the trade-off between label quality and model bias, as well as the limited adaptability arising from the reliance on manually set fixed thresholds. To address these challenges, we propose Bi-CoG, a plug-and-play self-training method that incorporates a deeply coupled pseudo-labeling strategy with error-aware dynamic filtering capabilities. Theoretical analysis and experimental results across 14 datasets in both open-world generalization and semi-supervised settings demonstrate that Bi-CoG can be seamlessly integrated with any pretrain-finetuning methodologies, consistently achieving performance improvements.
Acknowledgments
This research was supported by the Key Program of Jiangsu Science Foundation (BK20243012), the National Science Foundation of China (62306133), and the “111 Center” (No. B26023).
Appendix A Supplementary Material
A.1 Detailed Theoretical Analysis
A.1.1 Proof of Lemma 1
@@
When the aggregated training data contains noisy samples, let $\mathcal{F}$ denote the hypothesis space, $\eta$ the noise ratio in the training set ($\eta<0.5$), $\epsilon>0$ the target error rate, and $\delta<1$ the confidence parameter. According to PAC learning theory Angluin and Laird [1988]; Zhou and Li [2005], if the number of training samples $m$ satisfies
@@ A.1.2 Proof of Theorem 1
ImageNet-100 Natural picture 100 130,000 5,000
@@ A.2 Implementation Details
We evaluate our method on 14 recognition datasets, comprising 11 commonly used benchmarks for evaluating vision-language models (VLMs) and 3 classical benchmarks for semi-supervised learning. These datasets include CIFAR-10 and CIFAR-100 Krizhevsky et al. [2009], ImageNet and ImageNet-100 Deng et al. [2009], Caltech101 Fei-Fei et al. [2004], OxfordPets Parkhi et al. [2012], StanfordCars Krause et al. [2013], Food101 Bossard et al. [2014], Flowers102 Nilsback and Zisserman [2008], FGVCAircraft Maji et al. [2013], SUN397 Xiao et al. [2010], UCF101 Soomro et al. [2012], DTD Cimpoi et al. [2014], and EuroSAT Helber et al. [2019].
For ImageNet, Caltech101, OxfordPets, StanfordCars, Flowers102, FGVCAircraft, SUN397, UCF101, DTD, and EuroSAT, we follow the dataset splits used in CoOp Zhou et al. [2022a]. For CIFAR-10, CIFAR-100, and ImageNet-100, we adopt the standard semi-supervised splits used in FixMatch Sohn et al. [2020]. Detailed statistics of these datasets are provided in Table 7. Specifically, for the open-world generalization experiment, we follow the standard base-to-novel settingZhou et al. [2022a], where base and novel classes are split evenly. We sample 16 shots per class as labeled data, and treat all remaining training samples as unlabeled. Due to computational constraints, for ImageNet we randomly sample 5% of the unlabeled set for training.
@@
As a plug-and-play method, Bi-CoG can be seamlessly integrated into any pretrain–finetune paradigm. In our experiments, we adopt CoOp Zhou et al. [2022b], MaPLe Khattak et al. [2023a], and PromptSRC Khattak et al. [2023b] as the base methods, keeping all hyperparameters identical to those reported in their original papers. Unless otherwise specified, we set the number of models to $K=3$ and the scaling factor to $\alpha=1$. In the dynamic self-training stage, each model is first warmed up for $\zeta$ epochs using a small set of labeled data. Then, we apply our three-stage pseudo-label selection strategy to obtain pseudo-labels with high accuracy and low model bias, which are further used for self-training over $\kappa$ additional epochs. The detailed parameters are provided in Table 7. During the intra-model consistency stage, we apply weak and strong augmentations to each sample, with the specific augmentation strategies summarized in Table 6. We conduct all experiments on a single NVIDIA A800 GPU.
@@
A.3 Additional Experiments
Method Accuracy (%)
Self-training 95.02
FlexMatch (RA) 93.40 $\pm$ 0.05
CoMatch (RA) 93.39 $\pm$ 0.04
FixMatch (RA) 93.23 $\pm$ 0.14
Zero-shot CLIP 14.60
PromptSRC 59.97 $\pm$ 2.09
PromptSRC + Bi-CoG 61.03 $\pm$ 2.03
Table 8: Performance comparison of different methods on the ISIC2018 medical image dataset.
To further validate the effectiveness of Bi-CoG in domain-specific scenarios with limited labeled data, we conduct experiments on the ISIC2018 dataset Codella et al. [2018], which focuses on skin lesion diagnosis and comprises annotated images across seven distinct categories of skin diseases.
For conventional SSL methods (Self-training, FlexMatch Zhang et al. [2021], CoMatch Li et al. [2021a], and FixMatch Sohn et al. [2020]), we directly report the results from Zenk et al. [2022] to ensure consistency with established experimental benchmarks. For Bi-CoG, we adopt PromptSRC Khattak et al. [2023b] as the base model—an exemplar prompt tuning method tailored for vision-language models (VLMs). For statistical robustness, all results for Bi-CoG and PromptSRC are averaged over three runs with distinct random seeds. As illustrated in Table 8, Zero-shot CLIP achieves a mere 14.60% accuracy on this dataset. This poor performance stems from the paucity of medical image samples in CLIP’s pre-training corpus, leading to a severe mismatch between its generic visual features and the specialized requirements of skin lesion diagnosis. After prompt-based fine-tuning, PromptSRC significantly boosts the accuracy to 59.97%, demonstrating the effectiveness of task-specific prompt tuning in adapting VLMs to medical domains. When integrated with Bi-CoG, PromptSRC further yields 1.06% improvement to 61.03%, verifying that Bi-CoG’s adaptive pseudo-labeling strategy can effectively refine model performance even in specialized medical scenarios.
Notably, conventional SSL methods exhibit a substantial performance lead over fine-tuned VLMs on this dataset, with Self-training achieving the highest accuracy of 95.02%. This stark performance gap highlights a critical insight: constrained by the inherent limitations of their generic pre-trained features, VLMs still lag far behind end-to-end SSL algorithms in data-scarce medical imaging domains. Thus, for scenarios where labeled data is extremely limited and domain specificity is strong, it is more advisable to adopt traditional end-to-end SSL approaches rather than relying solely on pre-trained VLMs for downstream adaptation.